OpenStudy (anonymous):

A 25 degree angle has an opposite leg 6 cm long. How long is the adjacent leg?

OpenStudy (primeralph):

okay.....tan = O/A, tan 25deg = 6/A. Find A

OpenStudy (anonymous):

so i multiply the 6 and 25?

OpenStudy (primeralph):

chill......

OpenStudy (primeralph):